# Wasserstein Spatial Depth

### v1

@@ A.3 Vanishing at infinity
where we used the fact that $T_{Q_{n},P}(\mathbf{X}_{n})\sim P$ for all $n\in\mathbb{N}$. Since, $\sup_{n\in\mathbb{N}}\int\|\mathbf{x}\|^{2}{\rm d}Q_{n}(\mathbf{x})$ and $\sup_{P\in\mathcal{V}_{\beta}}\int\|\mathbf{x}\|^{2}{\rm d}P(\mathbf{x})$ are bounded, the previous display, (24) and (25) imply that the second term of (23) is uniformly integrable. Since $P$ and $P^{\prime}$ are exchangeable, the same holds for the third term. The uniform integrability of the last one follows directly from (24).
@@
From [43, Corollary 5.23], for every $\epsilon>0$, it holds that $Q(|T_{Q,P_{n}}-T_{Q,P}|\geq\epsilon)\to 0$. As $\|T_{Q,P_{n}}-T_{Q,P}\|_{L^{2}(Q)}$ is uniformly bounded, the sequence $\{T_{Q,P_{n}}-T_{Q,P}\}_{n\in\mathbb{N}}$ is compact w.r.t. the weak topology of $L^{2}(Q)$ by the Banach-Alaoglu–Bourbaki theorem (cf. [8, Theorem 3.16]). Therefore, for each subsequence $\{T_{Q,P_{n_{k}}}-T_{Q,P}\}_{k\in\mathbb{N}}$ there exists a further subsequence $\{T_{Q,P_{n_{k_{\ell}}}}-T_{Q,P}\}_{\ell\in\mathbb{N}}$ such that
@@
for some $L\in L^{2}(Q)$ and all $h\in L^{2}(Q)$. We prove now that $L=0$, irrespective of the subsequences. To improve readability, we write $\{T_{Q,P_{n}}-T_{Q,P}\}_{n\in\mathbb{N}}$ instead of $\{T_{Q,P_{n_{k_{\ell}}}}-T_{Q,P}\}_{\ell\in\mathbb{N}}$. Since $Q(|T_{Q,P_{n}}-T_{Q,P}|\geq\epsilon)\to 0$ and
@@
Arguing as in Section A.1 we get that
$\left\|\int_{V^{c}}\frac{I-T_{Q,P}}{\mathcal{W}_{2}(P,Q)}\mathrm{d}{\bf P}_{n}(P)\right\|_{L^{2}(Q)}+\left\|\int_{V^{c}}\frac{I-T_{Q,P}}{\mathcal{W}_{2}(P,Q)}\mathrm{d}{\bf P}_{n}(P)\right\|_{L^{2}(Q)}\leq{\bf P}(V^{c})+{\bf P}_{n}(V^{c})\leq\epsilon.$
@@
is compact (note that $V$ is compact in $\mathcal{P}_{2}(\mathbb{R}^{d})$ and $\mathcal{P}_{2}(\mathbb{R}^{d})\backslash\{Q\}\ni P\mapsto\frac{I-T_{Q,P}}{\mathcal{W}_{2}(P,Q)}$ is continuous, see Lemma 5.5), its closed convex hull, namely $C$, is compact as well. Since $\int_{V}\frac{I-T_{Q,P}}{\mathcal{W}_{2}(P,Q)}\mathrm{d}{\bf P}_{n}$ lies in $C$ for all $n\in\mathbb{N}$, the convergence of $\int_{V}\frac{I-T_{Q,P}}{\mathcal{W}_{2}(P,Q)}\mathrm{d}({\bf P}_{n}-{\bf P})(P)$ towards zero holds in the strong topology of $L^{2}(Q)$. We have proven that $A\leq\epsilon$ for $n$ big enough. Since $\epsilon$ was arbitrarily chosen, the result follows. ∎
@@
The latter tends to zero by Glivenko–Cantelli theorem and the fact that

### v2

@@ A.3 Vanishing at infinity
where we used the fact that $T_{Q_{n},P}(\mathbf{X}_{n})\sim P$ for all $n\in\mathbb{N}$. Since, $\sup_{n\in\mathbb{N}}\int\|\mathbf{x}\|^{2}{\rm d}Q_{n}(\mathbf{x})$ and $\sup_{P\in\mathcal{V}_{\beta}}\int\|\mathbf{x}\|^{2}{\rm d}P(\mathbf{x})$ are bounded, the previous display, (C.6) and (C.7) imply that the second term of (C.5) is uniformly integrable. Since $P$ and $P^{\prime}$ are exchangeable, the same holds for the third term. The uniform integrability of the last one follows directly from (C.6). ∎
@@
From [58, Corollary 5.23], for every $\epsilon>0$, it holds that $Q(\|T_{Q,P_{n}}-T_{Q,P}\|\geq\epsilon)\to 0$. As $\|T_{Q,P_{n}}-T_{Q,P}\|_{L^{2}(Q)}$ is uniformly bounded, the sequence $\{T_{Q,P_{n}}-T_{Q,P}\}_{n\in\mathbb{N}}$ is compact w.r.t. the weak topology of $L^{2}(Q)$ by the Banach-Alaoglu–Bourbaki theorem (cf. [10, Theorem 3.16]). Therefore, for each subsequence $\{T_{Q,P_{n_{k}}}-T_{Q,P}\}_{k\in\mathbb{N}}$ there exists a further subsequence $\{T_{Q,P_{n_{k_{\ell}}}}-T_{Q,P}\}_{\ell\in\mathbb{N}}$ such that
@@
for some $L\in L^{2}(Q)$ and all $h\in L^{2}(Q)$. We prove now that $L=0$, irrespective of the subsequences. To improve readability, we write $\{T_{Q,P_{n}}-T_{Q,P}\}_{n\in\mathbb{N}}$ instead of $\{T_{Q,P_{n_{k_{\ell}}}}-T_{Q,P}\}_{\ell\in\mathbb{N}}$. Since $Q(\|T_{Q,P_{n}}-T_{Q,P}\|\geq\epsilon)\to 0$ and
@@
Arguing as in Section A.1 we get that, for $n\geq n_{0}$,
$\left\|\int_{V^{c}}\frac{I-T_{Q,P}}{\mathcal{W}_{2}(P,Q)}\mathrm{d}{\bf P}(P)\right\|_{L^{2}(Q)}+\left\|\int_{V^{c}}\frac{I-T_{Q,P}}{\mathcal{W}_{2}(P,Q)}\mathrm{d}{\bf P}_{n}(P)\right\|_{L^{2}(Q)}\leq{\bf P}(V^{c})+{\bf P}_{n}(V^{c})\leq\epsilon.$
@@
is compact (note that $V$ is compact in $\mathcal{P}_{2}(\mathbb{R}^{d})$ and $\mathcal{P}_{2}(\mathbb{R}^{d})\backslash\{Q\}\ni P\mapsto\frac{I-T_{Q,P}}{\mathcal{W}_{2}(P,Q)}$ is continuous, see Lemma 5.5), its closed convex hull, namely $C$, is compact as well. Since $\int_{V}\frac{I-T_{Q,P}}{\mathcal{W}_{2}(P,Q)}\mathrm{d}{\bf P}_{n}$ lies in $C$ for all $n\in\mathbb{N}$, the convergence of $\int_{V}\frac{I-T_{Q,P}}{\mathcal{W}_{2}(P,Q)}\mathrm{d}({\bf P}_{n}-{\bf P})(P)$ towards zero holds in the strong topology of $L^{2}(Q)$. We have proven that $A\leq 2\epsilon$ for $n$ big enough. Since $\epsilon$ was arbitrarily chosen, the result follows. ∎
@@
The latter tends to zero by Glivenko–Cantelli theorem and the fact that, conditionally to $P_{1}$,
@@
{funding}
François Bachoc was supported by the Project GAP (ANR-21-CE40-0007) of the French National Research Agency (ANR) and by the Chair UQPhysAI of the Toulouse ANITI AI Cluster.